OpenStudy (anonymous):

I need to solve \[\sin x \cos x = \frac{ \sqrt{2} }{ 4 }\] Where do I begin and where do I end?

OpenStudy (cwrw238):

sin x cos x = 0.5 sin 2x 0.5 sin 2x = sqrt2/4 2 sin2x = sqrt2 sin 2x = sqrt2/2 now its easy - just look up the angle whose sine is sqrt2/2 on you calculator to get the value of 2x then divide by 2